# Dual-Stage Cross-Modal Network with Dynamic Feature Fusion for Emotional Mimicry Intensity Estimation

### v1

@@ Abstract
Emotional Mimicry Intensity (EMI) estimation serves as a critical technology for understanding human social behavior and enhancing human-computer interaction experiences, where the core challenge lies in dynamic correlation modeling and robust fusion of multimodal temporal signals. To address the limitations of existing methods in insufficient exploitation of modal synergistic effects, noise sensitivity, and limited fine-grained alignment capabilities, this paper proposes a dual-stage cross-modal alignment framework. First, we construct vision-text and audio-text contrastive learning networks based on an improved CLIP architecture, achieving preliminary alignment in the feature space through modality-decoupled pre-training. Subsequently, we design a temporal-aware dynamic fusion module that combines Temporal Convolutional Networks (TCN) and gated bidirectional LSTM to respectively capture the macro-evolution patterns of facial expressions and local dynamics of acoustic features. Innovatively, we introduce a quality-guided modality fusion strategy that enables modality compensation under occlusion and noisy scenarios through differentiable weight allocation. Experimental results on the Hume-Vidmimic2 dataset demonstrate that our method achieves an average Pearson correlation coefficient of 0.35 across six emotion dimensions, outperforming the best baseline by 40%. Ablation studies further validate the effectiveness of the dual-stage training strategy and dynamic fusion mechanism, providing a novel technical pathway for fine-grained emotion analysis in open environments.
@@ 1 Introduction
Emotional mimicry stands as one of the core behaviors in human social interactions, where individuals unconsciously replicate others’ facial expressions, vocal tones, or body movements to rapidly establish emotional resonance and strengthen social bonds. This psychological mechanism not only holds significant implications for understanding human social behavior but also provides natural inspiration for affective computing technologies in artificial intelligence. With the rapid development of applications such as virtual assistants, emotional robots, and mental health assessment, enabling machines to accurately perceive and quantify dynamic changes in human emotions has become a central challenge in human-computer interaction research. Among these, automatic estimation of Emotional Mimicry Intensity (EMI), capable of capturing the continuity and nuanced variations in emotional transmission, has gradually emerged as a cutting-edge direction in affective computing.
However, existing methods still face significant bottlenecks in addressing EMI estimation tasks. First, emotional mimicry is inherently a multimodal temporal process – subtle changes in visual expressions and dynamic fluctuations in vocal prosody often intertwine nonlinearly. Current research predominantly employs unimodal analysis (e.g., facial video only) or simplistic feature concatenation strategies, resulting in underutilized inter-modal synergy. For instance, baseline methods using ViT+GRU for visual information achieve merely 0.09 average Pearson correlation coefficient, indicating that unimodal visual models struggle to capture temporal evolution patterns of emotional intensity. Second, fine-grained alignment of emotional intensity requires cross-modal semantic understanding capabilities. Although Wav2Vec2 audio features combined with linear layers attain a relatively higher performance of 0.24, their unimodal nature limits robustness in complex scenarios (e.g., strong noise, occlusion). Furthermore, existing fusion strategies (e.g., simple average fusion in baselines with only 0.25 improvement) fail to account for differential information contributions across modalities at varying emotional phases, leaving multimodal potential underdeveloped.
@@
Through systematic experiments on the Hume-Vidmimic2 dataset, this paper validates the effectiveness of the proposed method. Compared to baseline models, our framework achieves a breakthrough performance with an average Pearson correlation coefficient of 0.35 across six emotional dimensions, demonstrating 40% improvement over the best baseline (0.25).
@@ 2.1 Multimodal Emotion Recognition Techniques
Multimodal emotion recognition enhances the robustness of affective analysis by fusing complementary information from visual, audio, and textual modalities. Early work such as Dynamic Fusion Network (DFN) and Tensor Fusion Network (TFN)[20, 21], which dynamically weights features through inter-modal gating mechanisms, yet its static weight allocation struggles to adapt to temporal variations in emotional intensity. Recent studies focus on cross-modal interaction modeling: Tsai et al. (2019) [9] designed a Memory-Augmented Cross-modal Attention (MULT) that utilizes external memory banks to store modality-shared semantics but fails to explicitly address semantic discrepancies caused by modal heterogeneity. Hazarika et al. (2020) [10] introduced cross-modal contrastive loss to narrow inter-modal distribution gaps via adversarial training, though their single-stage alignment strategy cannot handle hierarchical correlations in multimodal data. Compared to these approaches, our method decouples cross-modal semantic alignment through a dual-stage contrastive alignment mechanism, effectively mitigating inter-modal information redundancy and conflicts.
2.2 Temporal Modeling and Positional Encoding
The temporal dynamics of emotional mimicry intensity demand models capable of capturing both long-range dependencies and local contextual patterns. For visual temporal modeling, Lea et al. (2017) [11] proposed Temporal Convolutional Networks (TCN), which capture long-range dependencies between video frames through stacked dilated convolutions. However, their fixed receptive fields struggle to adapt to the non-stationary characteristics of emotional intensity variations. In audio temporal modeling, Hershey et al. (2021) [12] extracted spectral features via Wav2Vec2.0’s self-supervised pretraining but neglected explicit modeling of phoneme-emotion intensity mapping relationships. Regarding positional encoding, Vaswani et al. (2017) [13] introduced absolute positional encoding in Transformers, yet its adaptability for multimodal temporal alignment remains limited. Huang et al. (2020) [14] developed Modality-Adaptive Relative Positional Encoding (MARPE), which distinguishes video, audio, and textual characteristics through learnable parameters to facilitate differential modality fusion. Nevertheless, this approach overlooks the impact of modality quality fluctuations on fusion effectiveness. To address these limitations, we design a modality-aware dynamic fusion mechanism that automatically reduces the contribution weights of impaired modalities when quality discrepancies exist in multimodal signals (e.g., facial occlusion or environmental noise). Our framework combines TCN and bidirectional LSTM to respectively model local-global temporal dependencies in visual and audio modalities, while employing segment-wise average pooling to suppress interference from noisy frames.
@@ 2.3 Cross-Modal Alignment Techniques
Cross-modal alignment aims to establish shared semantic spaces for multimodal data. Radford et al. (2021) [1] proposed the CLIP model, which achieves image-text semantic alignment through contrastive learning, yet without extending to the audio modality. Wu et al. (2022) [15] constructed an audio-text joint embedding space (CLAP) based on the CLIP architecture, but their single-stage alignment strategy fails to balance multimodal semantic consistency. Recent work like Alayrac et al. (2022) [16] introduced the Flamingo model, achieving fine-grained image-text alignment through cross-attention, though its computational complexity limits practical deployment in real-time tasks. In affective computing, Yang et al. (2023) [17] aligned visual-audio-text features using contrastive learning but failed to explicitly model cross-modal uncertainty. This paper proposes a dual-stage alignment framework: The first stage independently aligns image-text and audio-text pairs through CLIP and CLAP architectures, while the second stage introduces a modality mapping layer to optimize triple-modality joint semantics, significantly enhancing fine-grained consistency in emotional intensity estimation.
@@ 3.1 Multimodal Data Preprocessing
Text Modality: The text processing pipeline begins with speech-to-text conversion using the Whisper model , which employs an end-to-end deep learning architecture optimized for speech recognition. Unlike conventional ASR systems, Whisper innovatively adopts a multi-task training paradigm, simultaneously optimizing speech recognition, language identification, and voice activity detection across 1.28 million hours of diverse audio data. This approach significantly improves robustness in noisy environments and multilingual adaptability. Specifically, we utilize a pruned and fine-tuned version, Whisper-large-v3-turbo , for high-accuracy transcription to obtain semantically complete utterances $s$.
Transcribed texts are encoded through the pretrained BERT model  using its 12-layer Transformer architecture to extract 768-dimensional context-aware feature vectors $f_{s}$. As a foundational NLP model, BERT’s dual pretraining objectives (masked language modeling and next-sentence prediction) enable effective capture of deep semantic relationships, providing reliable linguistic representations for our framework.
@@ 3.3 Dual-Stage Cross-Modal Alignment
Stage I: Modality-Text Contrastive Alignment Building upon the CLIP architecture [radford2021learning], we establish dual-path contrastive learning networks for vision-text and audio-text pairs. For visual modality, given video frame sequence $V=\{v_{t}\}_{t=1}^{T}$, frame-wise features $f_{v}^{t}\in\mathbb{R}^{d_{v}}$ are extracted through a ViT encoder, with segment-wise average pooling yielding video-level global features $\bar{f}_{v}$. Simultaneously, textual modality employs a pretrained BERT encoder to map transcribed speech $S$ into text features $f_{s}\in\mathbb{R}^{d_{s}}$. The vision-text contrastive loss is defined as:
@@
where $\tau$ denotes the temperature coefficient and $N$ the batch size. For audio modality, the CLAP architecture  processes acoustic signals $A$ through Wav2Vec2.0  to extract spectral features $f_{a}\in\mathbb{R}^{d_{a}}$, with analogous audio-text contrastive loss $\mathcal{L}_{a2t}$ following Equation (1).
@@ 4.1 Datasets and Experimental Setup
Experiments are conducted on the Hume-Vidmimic2 multimodal affective dataset, which contains 15,000 video samples spanning six emotional dimensions (”Admiration”, ”Amusement”, ”Determination”, ”Empathic Pain”, ”Excitement”, ”Joy”). Each sample’s emotional intensity score, ranging within $[0,1]$, is annotated by professional annotation teams following standardized protocols. The dataset is partitioned into training, validation, and test sets. Statistical details of each split are provided in Table 1.
@@ 5.2 Multimodal Fusion Effectiveness
  • The visual-textual pair achieves $\rho_{val}$=0.2934, surpassing audio-visual fusion (0.2846) by 3.1%
  • The audio-textual pair demonstrates strong synergy with $\rho_{val}$=0.3518, highlighting text modality’s complementary role
  • The full multimodal integration achieves state-of-the-art performance ($\rho_{val}$=0.3626), showing 27.5% improvement over best unimodal results
The progressive performance gains confirm our hypothesis about cross-modal complementarity. Notably, the textual modality contributes disproportionately to system performance, suggesting linguistic context plays a crucial role in disambiguating emotional mimicry patterns. Our dynamic fusion mechanism successfully preserves modality-specific information while enabling effective cross-modal interaction, as evidenced by the super-additive performance improvement in three-modality fusion.
@@
ViT Wav2Vec2.0 0.2846
ViT BERT 0.2934
Wav2Vec2.0 BERT 0.3518
ViT Wav2Vec2.0 BERT 0.3626
@@ 6 Conclusion
This paper presents our solution to the Emotional Mimicry Intensity (EMI) estimation challenge in the 8th Affective Behavior Analysis in-the-wild (ABAW) Competition. By introducing a dual-stage training paradigm and quality-aware dynamic modality fusion mechanism, our method achieves enhanced robustness in EMI estimation while maintaining tight three-modality alignment. Systematic validation on the Hume-Vidmimic2 dataset demonstrates that our approach attains an average Pearson correlation coefficient of 0.35 across six emotional dimensions, representing a 40% improvement over state-of-the-art baselines.

### v2

@@ Abstract
Emotional Mimicry Intensity (EMI) estimation serves as a critical technology for understanding human social behavior and enhancing human-computer interaction experiences, where the core challenge lies in dynamic correlation modeling and robust fusion of multimodal temporal signals. To address the limitations of existing methods in insufficient exploitation of modal synergistic effects, noise sensitivity, and limited fine-grained alignment capabilities, this paper proposes a dual-stage cross-modal alignment framework. First, we construct vision-text and audio-text contrastive learning networks based on an improved CLIP architecture, achieving preliminary alignment in the feature space through modality-decoupled pre-training. Subsequently, we design a temporal-aware dynamic fusion module that combines Temporal Convolutional Networks (TCN) and gated bidirectional LSTM to respectively capture the macro-evolution patterns of facial expressions and local dynamics of acoustic features. Innovatively, we introduce a quality-guided modality fusion strategy that enables modality compensation under occlusion and noisy scenarios through differentiable weight allocation. Experimental results on the Hume-Vidmimic2 dataset demonstrate that our method achieves an average Pearson correlation coefficient of 0.51 across six emotion dimensions, providing a novel technical pathway for fine-grained emotion analysis in open environments.
@@ 1 Introduction
Emotional mimicry stands as one of the core behaviors in human social interactions, where individuals unconsciously replicate others’ facial expressions, vocal tones, or body movements to rapidly establish emotional resonance and strengthen social bonds. This psychological mechanism not only holds significant implications for understanding human social behavior but also provides natural inspiration for affective computing technologies in artificial intelligence. With the rapid development of applications such as virtual assistants, emotional robots, and mental health assessment, enabling machines to accurately perceive and quantify dynamic changes in human emotions has become a central challenge in human-computer interaction research[6, 5]. Among these, automatic estimation of Emotional Mimicry Intensity (EMI), capable of capturing the continuity and nuanced variations in emotional transmission, has gradually emerged as a cutting-edge direction in affective computing.
However, existing methods still face significant bottlenecks in addressing EMI estimation tasks. First, emotional mimicry is inherently a multimodal temporal process – subtle changes in visual expressions and dynamic fluctuations in vocal prosody often intertwine nonlinearly. Current research predominantly employs unimodal analysis (e.g., facial video only) or simplistic feature concatenation strategies, resulting in underutilized inter-modal synergy. For instance, baseline methods using ViT[7] for visual information achieve merely 0.09 average Pearson correlation coefficient, indicating that unimodal visual models struggle to capture temporal evolution patterns of emotional intensity. Second, fine-grained alignment of emotional intensity requires cross-modal semantic understanding capabilities. Although Wav2Vec2[2] audio features combined with linear layers attain a relatively higher performance of 0.24, their unimodal nature limits robustness in complex scenarios (e.g., strong noise, occlusion). Furthermore, existing fusion strategies fail to account for differential information contributions across modalities at varying emotional phases, leaving multimodal potential underdeveloped.
@@
Through systematic experiments on the Hume-Vidmimic2 dataset, this paper validates the effectiveness of the proposed method. Compared to baseline models, our framework achieves a breakthrough performance with an average Pearson correlation coefficient of 0.51 across six emotional dimensions, demonstrating 200% improvement over the best baseline (0.25).
@@ 2.1 Multimodal Emotion Recognition Techniques
Multimodal emotion recognition enhances the robustness of affective analysis by fusing complementary information from visual, audio, and textual modalities. Early work such as Dynamic Fusion Network (DFN) and Tensor Fusion Network (TFN)[33, 34], which dynamically weights features through inter-modal gating mechanisms, yet its static weight allocation struggles to adapt to temporal variations in emotional intensity. Recent studies focus on cross-modal interaction modeling: Multimodal Transformer (MulT)[31] attends to interactions between multimodal sequences across distinct time steps and latently adapt streams from one modality to another,but fails to explicitly address semantic discrepancies caused by modal heterogeneity. MISA [10] projects each modality into two distinct subspaces. The first subspace learns their commonalities and reduces inter-modal discrepancies, while the second subspace captures the modality-specific characteristics of each modality.It exhibits excellent generalization capabilities across diverse sentiment analysis tasks. Compared to these approaches, our framework combines TCN[27] and bidirectional LSTM[35] to respectively model local-global temporal dependencies in visual and audio modalities, while employing segment-wise average pooling and modality-aware dynamic fusion mechanism to suppress interference from noisy frames.
@@ 2.2 Cross-Modal Alignment Techniques
Cross-modal alignment aims to establish shared semantic spaces for multimodal data. CLIP[28] achieves image-text semantic alignment through contrastive learning, yet without extending to the audio modality. CLAP[32] constructed an audio-text joint embedding space based on the CLIP architecture, but their single-stage alignment strategy fails to balance multimodal semantic consistency. Recent work introduced the Flamingo model[1], achieving fine-grained image-text alignment through cross-attention, though its computational complexity limits practical deployment in real-time tasks. In affective computing, ImageBind[8] learns a joint embedding across six different modalities - images, text, audio, depth, thermal, and IMU data. It enables novel emergent applications ‘out-of-the-box’ including cross-modal retrieval, composing modalities with arithmetic, cross-modal detection and generation.This paper proposes a dual-stage alignment framework: The first stage independently aligns image-text and audio-text pairs through CLIP and CLAP architectures, while the second stage introduces a modality mapping layer to optimize triple-modality joint semantics, significantly enhancing fine-grained consistency in emotional intensity estimation.
@@ 3.1 Multimodal Data Preprocessing
Text Modality: The text processing pipeline begins with speech-to-text conversion using the Whisper model [29], which employs an end-to-end deep learning architecture optimized for speech recognition. Unlike conventional ASR systems, Whisper innovatively adopts a multi-task training paradigm, simultaneously optimizing speech recognition, language identification, and voice activity detection across 1.28 million hours of diverse audio data. This approach significantly improves robustness in noisy environments and multilingual adaptability. Specifically, we utilize a pruned and fine-tuned version, Whisper-large-v3-turbo, for high-accuracy transcription to obtain semantically complete utterances $s$.
Transcribed texts are encoded through the pretrained BERT model [4] using its 12-layer Transformer architecture to extract 768-dimensional context-aware feature vectors $f_{s}$. As a foundational NLP model, BERT’s dual pretraining objectives (masked language modeling and next-sentence prediction) enable effective capture of deep semantic relationships, providing reliable linguistic representations for our framework.
@@ 3.3 Dual-Stage Cross-Modal Alignment
Stage I: Modality-Text Contrastive Alignment Building upon the CLIP architecture [28], we establish dual-path contrastive learning networks for vision-text and audio-text pairs. For visual modality, given video frame sequence $V=\{v_{t}\}_{t=1}^{T}$, frame-wise features $f_{v}^{t}\in\mathbb{R}^{d_{v}}$ are extracted through a ViT encoder, with segment-wise average pooling yielding video-level global features $\bar{f}_{v}$. Simultaneously, textual modality employs a pretrained BERT encoder to map transcribed speech $S$ into text features $f_{s}\in\mathbb{R}^{d_{s}}$. The vision-text contrastive loss is defined as:
@@
where $\tau$ denotes the temperature coefficient and $N$ the batch size. For audio modality, the CLAP architecture [32] processes acoustic signals $A$ through Wav2Vec2.0 [2] to extract spectral features $f_{a}\in\mathbb{R}^{d_{a}}$, with analogous audio-text contrastive loss $\mathcal{L}_{a2t}$ following Equation (1).
@@ 3.5 Quality-Aware Dynamic Fusion
Recent study [9] conducted a systematic multimodal robustness analysis of the Hume-Vidmimic2 dataset. By constructing a hierarchical evaluation framework under visual quality degradation scenarios, they revealed the significant impact of Facial Visual Quality (FVQ) on multimodal affective computing models. Critical interference factors such as partial occlusions (e.g. hair occluding eye/eyebrow regions) and abnormal illumination fluctuations (side overexposure-induced facial highlight noise) were shown to increase error rates in facial Action Unit (AU) detection and further induce cross-modal representation misalignment phenomena.
@@ 4.1 Datasets and Experimental Setup
Experiments are conducted on the Hume-Vidmimic2 multimodal affective dataset, which contains 15,000 video samples spanning six emotional dimensions (”Admiration”, ”Amusement”, ”Determination”, ”Empathic Pain”, ”Excitement”, ”Joy”). Each sample’s emotional intensity score, ranging within $[0,1]$, is annotated by professional annotation teams following standardized protocols. The dataset is partitioned into training, validation, and test sets. Statistical details of each split are provided in Table  1.
@@ 5.2 Multimodal Fusion Effectiveness
  • The audio-visual pair achieves $\rho_{val}$=0.3746, surpassing visual-textual fusion (0.2834) by 32.1%
  • The audio-textual pair demonstrates strong synergy with $\rho_{val}$=0.4418, highlighting text modality’s complementary role
  • The full multimodal integration achieves the best performance ($\rho_{val}$=0.5126),
The progressive performance gains confirm our hypothesis about cross-modal complementarity. Notably, the textual modality contributes disproportionately to system performance, suggesting linguistic and audio context plays a crucial role in disambiguating emotional mimicry patterns. Our dynamic fusion mechanism successfully preserves modality-specific information while enabling effective cross-modal interaction, as evidenced by the super-additive performance improvement in three-modality fusion.
@@
ViT BERT 0.2834
ViT Wav2Vec2.0 0.3746
Wav2Vec2.0 BERT 0.4418
ViT Wav2Vec2.0 BERT 0.5126
@@ 6 Conclusion
This paper presents our solution to the Emotional Mimicry Intensity (EMI) estimation challenge in the 8th Affective Behavior Analysis in-the-wild (ABAW) Competition. By introducing a dual-stage training paradigm and quality-aware dynamic modality fusion mechanism, our method achieves enhanced robustness in EMI estimation while maintaining tight three-modality alignment. Systematic validation on the Hume-Vidmimic2 dataset demonstrates that our approach attains an average Pearson correlation coefficient of 0.51 across six emotional dimensions.Provide a new approach for emotional mimicry intensity estimation.